# MolecularGPT: Open Large Language Model (LLM) for Few-Shot Molecular Property Prediction

### v1

@@ 4.3 Tuning on Hybrid Instruction Set
To investigate the RQ2, we conduct experiments to study the effect of hybrid instruction tuning set as presented in Fig. 4 and 5.
Tuning on property descriptions without demonstrations can improve the zero-shot performance. As shown in the 0-shot_tuning in Fig. 45, the model performed satisfactorily on some tasks under zero-shot inference but poorly on many tasks under few-shot inference. We speculated that the zero-shot instruction set imparts some knowledge to LLaMA without significantly enhancing the model’s ICL ability.
Providing the model with rich retrieved demonstrations would significantly improve its ICL ability. To test this, we fine-tuned the model on a 4-shot instruction dataset, represented by the 4-shot_tuning in Fig. 4 and 5. The results indicate an improvement in the model’s ICL ability. However, the model’s zero-shot generalization remained subpar on many tasks. We surmise that the model may learn shortcuts from the label words of the reference molecules rather than extracting the true relationships between the molecular representations and their properties.
Figure 4: The performance of MolecularGPT on classifcation tasks tuning with different types of instruction datasets. We inference them with 0, 1, and 2-shot examples. (0&4-shot indicates hybrid of 00 and $4$-shot. 0-4-shot indicates mix of 0,1,2,3,4-shot. tuning_double indicates double the instruction set size.)
Figure 5: The performance of MolecularGPT on regrassion tasks tuning with different types of instruction datasets. The setting is same as in Fig. 4.
@@
Mixed-shot instruction sets are promising to optimize both zero-shot generalization and ICL abilities. We developed two mixed instruction datasets: a combined 0&4-shot and a comprehensive mix of $0,1,2,3,4$-shot (0-4-shot) instruction datasets. As shown in 0&4-shot_tuning and 0-4-shot_tuning in Fig. 4 and 5, models fine-tuned on mixed-shot instruction datasets demonstrate a significant performance improvement compared to those fine-tuned on 0-shot or 4-shot instruction sets. This trend is consistently observed across various tested scenarios, indicating that our model derives the most benefit from mixed-shot instruction sets.
Tuning on larger instruction set have exhibited superior performance across different tasks under both zero and few shot learning. Models trained with larger datasets have exhibited superior performance on multi functional tasks, as evidenced by the improvements from GPT-2 (Radford et al., 2019) to GPT3 (Brown et al., 2020) and LLaMA2 (Touvron et al., 2023) to LLaMA3 (Meta, 2024). To further enhance MolecularGPT, we double the size of the 0&4-shot instruction sets. The results represented by the 0&4-shot_tuning_double in Fig. 4 and 5 suggest that expanding the data scale enhances the model’s performance across various tasks either by zero-shot or few-shot learning.
@@ 4.4 Hyperparameter Sensitivity Analysis
To fully utilize the ICL ability of MolecularGPT, we now pay attention to the impact of number, order and diversity of in-context demonstrations to discuss the RQ3. The results are depicted in Fig 6.
@@
MolecularGPT gains significant enhancement with up to $2$ demonstrations, but the marginal benefit diminishes with additional retrieval molecules. We investigate the impact of the number (Ye et al., 2024) of retrieval demonstrations, ranging from 00 to $8$ examples based on similarity. The results indicate significant improvement when provided with up to $2$ examples on many datasets. However, the performance does not get further improvement with more retrieval molecules. We hypothesize that: 1) More noise will be introduced with the increase of examples that has lower similarity with the query. 2) The maximum input length of $512$ tokens with at most $4$ examples in instructions constrains the model’s capability while handling more examples.
@@
Ascending order of similarity for demonstrations is sub-optimal compared to descending order especially with more demonstrations. We arrange the demonstrations (Lu et al., 2022; Zhao et al., 2021) in a ascending order, placing the most similar examples at the end of k-shot instructions. The results in Fig. 6 show that the ascending order is sub-optimal comparing to descending order, especially with more demonstrations which may be constrained by the model’s long context capability. We also assume the model is more adaptable to reasoning with descending order by learning most related knowledge first.
Similar retrieved molecule demonstrations provides better performance than diverse demonstrations. To increase the diversity, we retrieve equal number of molecules from each category (Ma et al., 2024). When the same number of examples is provided within instructions, the retrieval approach based on similarity consistently outperforms the one based on diversity across all classification tasks as shown in Fig. 6. The similarity-based methodology tends to provide examples that align more coherently with the query molecules. In contrast, the diversity-based approach offers a mix of positive and negative examples, which potentially introduce noise and create ambiguity perplexing the language models.
@@ A.1 Details of datasets
We adhere to the dataset selections as outlined in GIMLET (Zhao et al., 2024). Moreover, considering the importance and extensive research in the field of quantum mechanical properties, we have included an additional two quantum mechanical properties: Highest Occupied Molecular Orbital(HOMO) and Lowest Unoccupied Molecular Orbital (LUMO), from the QM9 datasets (Ramakrishnan et al., 2014) as our instruction tuning datasets. To construct the instructions for these additional datasets, we employed the method in Mol-Instructions (Fang et al., 2023) and GIMLET (Zhao et al., 2024). Initially, we write a property description for each task according to Wikipedia and chemistry papers. Subsequently, we employ GPT-4.0 (OpenAI, 2023) to generate instructions based on these seed examples, resulting in various human question-framing styles instructions. The comprehensive list of tuning and downstream tasks are summarized in Tab. 3.
@@ C.2 The effect of instruction datasets
In the combined 0&4-shot methodology, we merge the 0-shot and 4-shot instruction datasets in an equal ratio of 0.5: 0.5. For the comprehensive 0-4 shot mix, we integrate the 0,1,2,3, and 4-shot instruction datasets in a ratio of 0.6: 0.1: 0.1: 0.1: 0.1. During these procedures, we ensure the absence of duplicate query molecules and maintain the scale of the datasets. For the doubled scale of 0&4-shot, we amalgamate the 0-shot and 4-shot instruction datasets in an equal proportion of 1: 1. The results of the zero- and few-shot inferences are presented in the following Tab. 5678 and  9.

### v2

@@ 4.3 Tuning on Hybrid Instruction Set
To investigate the RQ2, we conduct experiments to study the effect of hybrid instruction tuning set as presented in Fig. 3.
Tuning on property descriptions without demonstrations can improve the zero-shot performance. As shown in the 0-shot_tuning in Fig. 3, the model performed satisfactorily on some tasks under 0-shot inference but poorly on many tasks under 2-shot inference. In comparison with 2-shot inference, 0-shot inference even demonstrates an average improvement of 4% on classification tasks. We speculated that the zero-shot instruction set imparts some knowledge to LLaMA without significantly enhancing the model’s ICL ability.
Providing the model with rich retrieved demonstrations would significantly improve its ICL ability. To test this, we fine-tuned the model on a 4-shot instruction dataset, represented by the 4-shot_tuning in Fig. 3. The results indicate an improvement in the model’s ICL ability. In 2-shot inference setting, comparing with 0-shot_tuning, 4-shot_tuning shows an average 8.2% increase on classification tasks. However, the model’s zero-shot generalization remained subpar on many tasks. We surmise that the model may learn shortcuts from the label words of the reference molecules rather than extracting the true relationships between molecular representations and properties.
Figure 3: The performance of MolecularGPT on Classifcation (Cls) and Regression (Reg) tasks tuning with different types of instruction datasets. We inference them with 0-shot and 2-shot examples. (0&4-shot indicates hybrid of 00 and $4$-shot. 0-4-shot indicates mix of 0,1,2,3,4-shot. tuning_double indicates double the instruction set size.)
@@
Mixed-shot instruction sets are promising to optimize both zero-shot generalization and ICL abilities. We developed two mixed instruction datasets: a combined 0&4-shot and a comprehensive mix of $0,1,2,3,4$-shot (0-4-shot) instruction datasets. As shown in 0&4-shot_tuning and 0-4-shot_tuning in Fig. 3, models fine-tuned on mixed-shot instruction datasets demonstrate a significant performance improvement compared to those fine-tuned on 0-shot or 4-shot instruction sets. This trend is consistently observed across various tested scenarios, like BACE, HIV, Tox21, BBBP and CYP450, indicating that our model derives the most benefit from mixed-shot instruction sets.
Tuning on larger instruction set have exhibited superior performance across different tasks under both zero and few shot learning. Models trained with larger datasets have exhibited superior performance on multi functional tasks, as evidenced by the improvements from GPT-2 (Radford et al., 2019) to GPT3 (Brown et al., 2020) and LLaMA2 (Touvron et al., 2023) to LLaMA3 (Meta, 2024). To further enhance MolecularGPT, we double the size of the 0&4-shot instruction sets. The results represented by the 0&4-shot_tuning_double in Fig. 3 suggest that expanding the data scale enhances the model’s performance across various tasks, like HIV, MUV and CYP450, either by zero-shot or few-shot learning.
@@ 4.4 Hyperparameter Sensitivity Analysis
To fully utilize the ICL ability of MolecularGPT, we now pay attention to the impact of question prompts and the retrieval examples to discuss the RQ3. Specifically, for question prompts, we test the robustness of our model respons to different instructions, shown in Tab. 3. As for few-shot demonstrations, we discuss the impact of molecular fingerprints in Tab. 4, as well as the number, order and diversity of demonstrations in Fig 4.
The exceptional robustness of MolecularGPT is validated across different tasks. Given the diversity and flexibility of natural language, we aim to evaluate the robustness of MolecularGPT against various instruction prompts. Adhering to the downstream datasets used in GIMLET, which provides five distinct types of instructions. We calculate the standard deviation of the ROC-AUC or RMSE metrics derived from these five instruction datasets. Comparative results with GIMLET is presented in Tab. 3. It is evident that our model demonstrates superior robustness compared to GIMLET across most tasks, with an average standard deviation of 0.06 compared to GIMLET’s 0.08. This indicts the robustness of MolecularGPT that it genuinely comprehends complex instructions and can handle a range of property prediction tasks without requiring task-specific prompt designs.
The effectiveness of MolecularGPT with different fingerprints for retrieving examples. In both the 2-shot and 8-shot settings, our model demonstrates significant effectiveness when utilizing MACCS or MACCS-ECFP4 fingerprints for retrieving in-context examples, as illustrated in Tab. 4. The method based on MACCS-ECFP4 even outperforms our original MACCS-based method on 7 out of 10 tasks in the 8-shot setting. This indicates that our model exhibits strong structural awareness and generalization ability.
Table 3: Standard deviation for GIMLET and MolecularGPT in response to 5 types of instructions.
Classification (AUC-ROC) Regression (RMSE)
Method BACE HIV MUV Tox21 ToxCast BBBP CYP450 ESOL FreeSolv Lipo
GIMLET 0.024 0.034 0.017 0.009 0.005 0.013 0.012 0.020 0.532 0.082
MolecularGPT 0.009 0.018 0.012 0.008 0.002 0.045 0.002 0.007 0.498 0.002
Table 4: The performance of MolecularGPT when using different fingerprints for retrieving in-context examples.
Classification (AUC-ROC) Regression (RMSE)
Method Shot BACE HIV MUV Tox21 ToxCast BBBP CYP450 ESOL FreeSolv Lipo
MACCS 2-shot 0.7218 0.7204 0.6338 0.6573 0.5945 0.7260 0.8275 1.489 5.226 1.015
MACCS-ECFP4 0.7675 0.6996 0.6042 0.6749 0.5993 0.7298 0.8406 1.502 5.085 1.039
MACCS 8-shot 0.7331 0.6382 0.6469 0.6565 0.5985 0.6822 0.8228 1.433 5.033 1.028
MACCS-ECFP4 0.7926 0.6674 0.6109 0.6685 0.5985 0.6962 0.8445 1.512 5.017 1.003
@@
MolecularGPT gains significant enhancement with up to $2$ demonstrations, but the marginal benefit diminishes with additional retrieval molecules. We investigate the impact of the number (Ye et al., 2024) of retrieval demonstrations, ranging from 00 to $8$ examples based on similarity. The results indicate significant improvement when provided with up to $2$ examples cross all datasets. However, the performance does not get further improvement with more retrieval molecules. We hypothesize that: 1) More noise will be introduced with the increase of examples that has lower similarity with the query. 2) The maximum input length of $512$ tokens with at most $4$ examples in instructions constrains the model’s capability while handling more examples.
@@
Ascending order of similarity for demonstrations is sub-optimal compared to descending order especially with more demonstrations. We arrange the demonstrations (Lu et al., 2022; Zhao et al., 2021) in a ascending order, placing the most similar examples at the end of k-shot instructions. The results in Fig. 4 show that the ascending order is sub-optimal comparing to descending order, especially with more demonstrations which may be constrained by the model’s long context capability. In an 8-shot setting, the ascending order shows an average performance that is 0.9% lower than the descending order on classification tasks. We assume the model is more adaptable to reasoning with descending order by learning most related knowledge first.
[…]
@@ A.1 Details of datasets
We follow the dataset selection and splitting criteria specified in GIMLET (Zhao et al., 2024). As for downstream tasks, we divide the datasets in a ratio of 0.8:0.1:0.1 and report the results on the test sets. Moreover, considering the importance and extensive research in the field of quantum mechanical properties, we have included an additional two quantum mechanical properties: Highest Occupied Molecular Orbital(HOMO) and Lowest Unoccupied Molecular Orbital (LUMO), from the QM9 datasets (Ramakrishnan et al., 2014) as our instruction tuning datasets. To construct the instructions for these additional datasets, we employed the method in Mol-Instructions (Fang et al., 2023) and GIMLET (Zhao et al., 2024). Initially, we write a property description for each task according to Wikipedia and chemistry papers. Subsequently, we employ GPT-4.0 (OpenAI, 2023) to generate instructions based on these seed examples, resulting in various human question-framing styles instructions. The comprehensive list of tuning and downstream tasks are summarized in Tab. 5.
@@ C.2 The effect of instruction datasets
In the combined 0&4-shot methodology, we merge the 0-shot and 4-shot instruction datasets in an equal ratio of 0.5: 0.5. For the comprehensive 0-4 shot mix, we integrate the 0,1,2,3, and 4-shot instruction datasets in a ratio of 0.6: 0.1: 0.1: 0.1: 0.1. During these procedures, we ensure the absence of duplicate query molecules and maintain the scale of the datasets. For the doubled scale of 0&4-shot, we amalgamate the 0-shot and 4-shot instruction datasets in an equal proportion of 1: 1. The results of the zero- and few-shot inferences are presented in the following Tab. 78910 and  11.